# Tag&Tab: Pretraining Data Detection in Large Language Models Using Keyword-Based Membership Inference Attack

### v1

@@ 4 Evaluation
In this section, we provide a detailed evaluation to showcase the effectiveness of Tag&Tab.
@@ 4.1 Model Comparison
To evaluate Tag&Tab in comparison to other reference-free baseline detection methods, we examined various open-source LLMs, including LLaMA 1 (7B, 13B, 30B) [7], Pythia (160M, 1.4B, 2.8B, 6.9B, 12B) [42]. Additionally, we used GPT-3.5 Turbo [^1] https://platform.openai.com/docs/models/gpt-3-5-turbo (trained on data up to September 2021222https://learn.microsoft.com/en-us/azure/ai-services/openai/concepts/models), given the partial knowledge that it was trained on known books, as discussed in previous studies [38, 44]. Models like LLaMA 1 and Pythia are well-suited for MIA evaluation with their transparency regarding pretraining datasets. This contrasts with newer models such as LLaMA 2 and 3, which lack such transparency.
@@ 4.2 Dataset Comparison
The experiments were conducted on the BookMIA333www.huggingface.co/datasets/swj0419/BookMIA [38], The Pile444www.huggingface.co/datasets/monology/pile-uncopyrighted/viewer/default/validation [12], and MIMIR555www.huggingface.co/datasets/iamgroot42/mimir [27] datasets, meeting the requirement that MIA evaluation datasets should be as comprehensive and diverse as possible [27], covering a wide range of text types while maintaining consistency in the distribution of text between the training and test sets.
The Pile is a collection of diverse textual sources designed to train and evaluate LLMs using open-source data, from which we utilized 10,000 samples each for training and testing from each domain. BookMIA focuses on evaluating MIAs with books known to have been memorized by GPT-3.5 Turbo, alongside newly published books from 2023, using 5,000 samples each for the member and non-member sets. It is important to note that 34 of the 50 ’member’ books in BookMIA overlap with the Gutenberg dataset, which is also known to have been part of the training corpora for models such as Pythia and LLaMA 1. MIMIR is a dataset built from The Pile, designed to evaluate memorization in LLMs. It contains data known to have been used in training across all Pythia model sizes, offering a unified benchmark for assessing membership inference. In our evaluation, we utilized around 2,000 samples per domain, combining approximately 1,000 ’member’ and 1,000 ’non-member’ samples.
It is important to note that we opted not to evaluate our method on the WikiMIA dataset666www.huggingface.co/datasets/swj0419/WikiMIA [38], as recent publications (e.g., [24]) questioned the reliability of the data, due to temporal shifts in writing styles and an insufficient number of samples.
@@ 4.3 Evaluation Approach
To assess our method’s performance, we followed a systematic process that begins with the input dataset. Each text file in the dataset is processed, with every sentence truncated to a maximum of 2,048 tokens to ensure a consistent input size. Sentences containing fewer than seven words are filtered out. We identify and save the top-K keywords for every sentence based on their entropy values. We documented the outcomes of selecting 1 to 10 keywords per sentence.
@@
Once the top-k keywords are selected, the entire text is fed into the target model, which outputs the probability distribution for each token. Then we average the log-likelihoods of all tokens in the keywords, conditioned on their preceding tokens, assessing the model’s familiarity with the entire keyword.
@@
Following the evaluation process suggested by Carlini et al.[26], we set a threshold value to determine attack performance, focusing on TPR at low FPR, denoted as T@F.
@@
In addition, we report the area under the ROC curve (AUC score), which enhances understanding of the detection. The AUC score quantifies the overall performance of a classification method by considering TPRs and FPRs at all classification thresholds. Since AUC offers a comprehensive, threshold-independent evaluation metric, we do not need to determine a specific threshold $\gamma$ for our method.
4.4 Comparison with Baseline Methods
To benchmark Tag&Tab’s performance, we compared it to the performance of SOTA reference-free methods for detecting pretraining data. The baseline methods included LOSS Attack, Zlib Attack, Neighbor Attack, Min-KProb, Min-K%++ Prob, Max-K%Prob, RECALL, DC-PDD. Comprehensive descriptions of these attacks can be found in the Related Work section.
5 Results
In this section, we present two distinct case studies that utilize Tag&Tab, each exploring a different aspect of pretraining data detection in LLMs. Each case study is evaluated using the following metrics: AUC and TPR at a low FPR of 5% (T@F=5%).
AUC T@F5 AUC T@F5 AUC T@F5 AUC T@F5 AUC T@F5 AUC T@F5 AUC T@F5
Neighbor 64.7 26.8 71.2 38.3 89.6 72.8 64.9 25.8 70.9 36.1 96.3 87.5 76.3 47.9
Loss 58.9 24.6 69.6 43.1 88.7 74.3 62.4 23.7 68.5 32.2 96.6 89.7 74.1 47.9
Zlib 52.5 21.9 66.8 41.6 88.7 74.1 54.6 18.7 61.1 25.4 95.6 88.1 69.9 44.9
Min-20.0% Prob 61.1 24.3 70.3 42.4 86.7 70.0 65.2 24.6 70.0 34.0 95.1 86.2 74.7 46.9
MinK++-20.0% Prob 59.7 23.0 67.7 37.6 78.0 60.2 59.1 19.6 56.2 20.1 95.1 86.2 69.3 41.1
Max-20.0% Prob 51.4 15.3 65.6 33.9 87.4 69.3 50.6 12.7 58.8 20.2 96.0 90.6 68.3 40.3
ReCaLL 57.8 22.3 69.8 41.9 84.1 64.0 66.2 29.1 71.7 36.8 73.7 50.3 70.5 41.1
DC-PDD 60.7 26.9 71.0 46.9 88.3 76.7 67.9 33.5 73.6 44.2 95.2 88.7 76.1 52.8
Ours (Tag&Tab K=4) 69.1 28.2 77.8 47.5 91.4 76.4 71.7 29.6 75.4 35.9 97.0 89.6 80.4 51.2
Ours (Tag&Tab K=10) 67.0 26.1 76.6 46.2 91.4 76.5 71.9 30.2 75.6 35.6 95.9 87.3 79.7 50.3
Table 1: Detection of data from the BookMIA dataset used in pretraining seven models using Tag&Tab and six baseline MIAs, evaluated in terms of the AUC and T@F=5%. The last two rows compare the Tag&Tab method when selecting four and ten keywords. The best results are highlighted in bold, and the second-best results are underscored.
5.1 Case Study 1: Detecting Specific Pretraining Data in LLMs
This case study focuses on the precision of detecting specific pretraining data in LLMs. To do so, we designed a well-engineered attack aimed at inferring whether copyrighted data was part of the pretraining of the examined LLMs. In contrast to Case Study 2, the data used for training the target model and the data identified as non-members here do not originate from the same database. Using the BookMIA dataset, we simulate the possibility of having knowledge of part of a model’s pretraining data, which can be used to infer specific text files suspected of being included in the pretraining data of the target model. For validation, non-members were collected from books that were released after the publication of the target model, ensuring that these texts could not be part of its pretraining data.
To determine the optimal number of keywords to select, we evaluate the results by selecting between 1 and 10 keywords from each sentence The results shown in Figure 2 demonstrate that the optimal number of keywords required to ensure effective detection depends on the model architecture. As observed, for different sizes of the LLaMa 1 models, the optimal number of keywords ranged from 2 to 3, while for the Pythia models and GPT-3.5 turbo, the optimal number of tagged keywords was 7. On average the best results across all models were obtained when $K=4$. resulting in an average AUC score of 80.4%.
Tag&Tab outperforms all other attacks in terms of the AUC score, achieving an average improvement in the AUC compared to the other attacks, ranging from 4.1%-12.1% for Tag&Tab K=4. The performance of Tag&Tab K=4 surpasses that of other attacks in terms of the T@F=5% on LLaMa-7b and LLaMa-13b models. However, for most other models, the DC-DPP attack achieves better results, with Tag&Tab consistently ranking second in performance. As the size of the tested LLM increases, the AUC scores of the MIAs also increase due to the model’s memorization capacity [29]. This can be seen in the results obtained for Tag&Tab which obtained very high AUC scores: (1) 91.4% on LLaMa-30b compared to 69.1% on LLaMa-7b, (2) 75.4% on Pythia-12b compared to 71.7% Pythia-6.9b, and (3) 97% on GPT-3.5 Turbo, which is the largest model tested.
Finally, we also investigated the impact of our method’s word selection stage, tagging, in which the highest $K$ entropy words are selected, compared to a random selection of tokens from the sentence. The results indicate that selecting the highest $K$ entropy words significantly improves performance across all models, achieving an average AUC score of 80.4%, compared to an average AUC score of 64.4% with a random selection of $K$ tokens.
Figure 2: AUC scores as a function of the number of tagged keywords for the examined models on the BookMIA dataset. Yellow dots indicate optimal performance: 2-3 keywords for LLaMa 1, 7 for Pythia and GPT-3.5 turbo, and 4 on average (AVG).
@@ 5.1 Case Study 1: Detecting Specific Pretraining Data in LLMs
Table 2: Comparison of AUC results for our method, Tag&Tab, against baseline methods on the MIMIR and The PILE benchmarks. The upper table presents the best results from Tag&Tab and baseline methods on four datasets from the MIMIR benchmark, while the lower table shows the best results on four datasets from The PILE benchmark. The best results for each dataset and model size are highlighted in bold, and the second-best AUC is underscored.
@@
5.2 Case Study 2: Detecting Various Types of Pretraining Data in LLMs
@@
This case study examines the Tag&Tab effectiveness and robustness in detecting various types of pretraining data in LLMs. We evaluate our method on various sizes of the Pythia model and compare its effectiveness against baseline attacks on 7 different types of textual data in the Pile dataset. Table 2 summarizes the results obtained when targeting five Pythia model sizes ranging from 160M to 12B parameters, tested with two configurations of tagged keywords: 4 and 10.
@@
Tag&Tab ($K=4$) consistently outperforms all baseline methods on average across the evaluated models, establishing itself as the most effective approach overall. Tag&Tab ($K=10$) ranks as the second-best method, demonstrating strong performance but falling short of the results achieved with $K=4$. Notably, Tag&Tab ($K=4$) achieves either the best or second-best results in the majority of tested types of textual daat. Even when it does not lead, its performance remains competitive, providing a robust alternative to the leading method.
@@
We can also see that while Tag&Tab ($K=10$) selects more keywords, thus increasing the number of probabilities considered for text membership inference, the results are consistently lower compared to Tag&Tab when $K=4$. This supports the hypothesis that selecting a smaller number of keywords allows the method to extract noise-free information from the model.
@@
Despite outperforming baseline membership inference attacks (MIAs), the AUC achieved by Tag&Tab can still be relatively low in certain cases, hovering around 55%. However, new research performed by Maini et al. [24] shows that aggregating multiple MIAs improves dataset membership inference accuracy, motivating the need for better attacks that achieve AUC over 50% for improved aggregated attack performance.
@@ 6 Conclusion
We present Tag&Tab, a novel black-box method for detecting pretraining data in LLMs. The method’s focus on the semantic and contextual relevance of words in the text enhances its detection capabilities. Tag&Tab was shown to outperform SOTA attacks and demonstrates consistently high performance across diverse types of textual data. In our comprehensive evaluation on 8 different types of textual data from three datasets (the Pile, MIMIR, and BookMIA datasets) and six different LLMs (LLaMa 1-7b, 13b, 30b, Pythia-160m, 1.4b, 2.4b, 6.9b, 12b, GPT-3.5 Turbo) of varying sizes and architectures. Our study confirms that the selection of high-entropy keywords improves membership inference attack results, further validating our approach. Future work could explore extending the Tag&Tab method by incorporating dimensions such as the context, and placement of keywords within the document. Additionally, developing new MIAs that leverage advanced NLP techniques to focus on the significance of words within the text could further enhance the ability to detect pertaining data in LLMs.

### v2

@@ 4 Evaluation
This section presents a detailed evaluation of Tag&Tab’s effectiveness. The experiments were conducted on a NVIDIA RTX 6000 GPU, running for nearly three days on all models and datasets. We used the default parameter settings of widely adopted libraries, including spaCy and NLTK.
@@ 4.1 Model Comparison
To compare Tag&Tab with other reference-free baseline detection methods, we examined various open-source LLMs, including LLaMA 1 (7B, 13B, 30B) [4], Pythia (160M, 1.4B, 2.8B, 6.9B, 12B) [34], and Qwen1.5-14B [39]. Additionally, we included GPT-3.5 Turbo [^1] https://platform.openai.com/docs/models/gpt-3-5-turbo (trained on data up to September 2021222https://learn.microsoft.com/en-us/azure/ai-services/openai/concepts/models), given partial knowledge of its training on known books, as discussed in previous studies [30, 40]. Our black-box assumption still holds because the OpenAI API exposes token-level log probabilities for this model. LLaMA 1 and Pythia are well-suited for MIA evaluation due to their transparency regarding pretraining datasets, unlike newer models such as LLaMA 2 and 3, which lack such transparency.
@@ 4.2 Dataset Comparison
The experiments were conducted on the BookMIA333www.huggingface.co/datasets/swj0419/BookMIA [30], Pile444www.huggingface.co/datasets/monology/pile-uncopyrighted/viewer/default/validation [9], and MIMIR555www.huggingface.co/datasets/iamgroot42/mimir [22] datasets, meeting the requirement that MIA evaluation datasets should be as comprehensive and diverse as possible [22], covering various types of text while maintaining a consistent distribution between training and test sets.
The Pile is a collection of diverse texts of different types from a wide range of sources, designed to train and evaluate LLMs using open-source data. In our experiments, for each domain we used 10,000 samples for training and an additional 10,000 samples from the same domain for testing. BookMIA is a dataset consisting of 100 books, with 100 text chunks of 512 words extracted from each, totaling 10,000 samples. The dataset is evenly split into member and non-member sets. The member set includes 5,000 samples from 50 books from the Books3 dataset [41], a collection of ebooks, most of which were published between 2000 and 2020. The non-member set comprises 5,000 samples from 50 new books with first editions in 2023 that could not have been in the training data of GPT-3.5 Turbo [40]. BookMIA is designed to evaluate MIAs by contrasting model behavior on seen and unseen content. For our evaluation on Pythia and LLaMA 1, we leveraged the fact that their training corpora, the Pile and Books3 respectively, have significant overlap with the Gutenberg dataset. The BookMIA member set, sourced from Books3, contains 50 books, 34 of which are also present in Gutenberg. To construct a high-confidence member set, we focused our evaluation on this specific overlap, filtering out the 16 member books from BookMIA that are not part of the Gutenberg corpus. This ensures that our test set consists of books with the strongest possible evidence of inclusion in the pretraining data, providing a more controlled experimental setup. We note that cutoff-based datasets such as BookMIA have been critiqued as unsuitable for strict benchmarking of MIAs due to temporal distribution shifts [42, 17]. Accordingly, we do not treat BookMIA as a primary benchmark but rather as an illustrative case study simulating a practical audit scenario. MIMIR is a dataset built from the Pile, designed to evaluate memorization in LLMs. MIMIR is a post-processed subset of the Pile, intended to reduce distributional artifacts such as temporal drift and stylistic shifts. It contains data known to have been used in training all Pythia model sizes, offering a unified benchmark for assessing membership inference. In our evaluation, we utilized around 2,000 samples per domain (approximately 1,000 member and 1,000 non-member samples).
The PatentMIA [^6] https://github.com/zhang-wei-chao/DC-PDD/tree/main [33] corpus was used in a dedicated non-Latin evaluation on the Chinese language, using the Qwen1.5-14B model. The setup and results are discussed in Appendix A.4.
It is important to note that we opted not to evaluate our method on the WikiMIA dataset,777www.huggingface.co/datasets/swj0419/WikiMIA [30] as recent publications (e.g., [17]) questioned the reliability of the data due to temporal shifts in writing styles and an insufficient number of samples.
@@ 4.3 Evaluation Approach
To assess our method’s performance, we followed a systematic process that begins with the input dataset. Each text file in the dataset is processed, with every document truncated to a maximum of 2,048 tokens to ensure a consistent input size. Sentences with fewer than seven words are excluded. We identify and save the K keywords for every sentence. We documented the outcomes of selecting 1 to 10 entropy keywords per sentence (before adding the NER words).
@@
After selecting the K keywords, the entire text is processed by the target model, which outputs probability distributions for each token. Then we average the log-likelihoods of all tokens in the keywords, conditioned on their preceding tokens, assessing the model’s familiarity with the entire keyword.
@@
Following Carlini et al.’s evaluation process [19], we set a threshold to assess attack performance, focusing on TPR at low FPR, denoted as T@F.
@@
We also report the area under the ROC curve (AUC score) to provide a clearer measure of detection performance. The AUC score quantifies the overall performance of a classification method by considering TPRs and FPRs at all classification thresholds. Since AUC offers a comprehensive, threshold-independent evaluation metric, we do not need to determine a specific threshold $\gamma$ for our method.
To simulate a real-world application, Appendix A.3 details how we pick a working threshold $\gamma$ when only non–member data from the same domain are available. The appendix further shows that applying this book-derived threshold to a different domain (mathematics) degrades performance, so $\gamma$ must be recalibrated for every model–dataset pair.
@@
Table 1: Comparison of AUC results for Tag&Tab and baseline methods on the MIMIR and Pile benchmarks. The upper table presents the best results from Tag&Tab and baseline methods across four MIMIR datasets, while the lower table shows the best results for three Pile datasets. The best results for each dataset and model size are highlighted in bold, and the second-best AUC is underscored.
@@
5 Results
This section presents the results of two case studies using Tag&Tab, each examining a different aspect of pretraining data detection in LLMs. Each case study is evaluated using the following metrics: AUC and TPR at a low FPR of 5% (T@F=5%).
Throughout this section, we report the results for Tag&Tab using $K=4$ and $K=10$, as these values yielded the most consistent and high-performing outcomes across models and datasets. This choice is supported by the results presented in Figure 2, which shows that Tag&Tab performs robustly for $K$ values between 4 and 10 for most of the models, with minimal performance variation—making the method resilient to non-optimal keyword selections.
Case Study 1 serves as the primary benchmark evaluation, while Case Study 2 is included to illustrate a more practical, engineered audit scenario closer to real-world use.
The reported results are based on a single run, as we observed minimal variation across multiple runs.
@@ 5 Results
5.1 Case Study 1: Detecting Various Types of Pretraining Data in LLMs
@@
This case study evaluates Tag&Tab’s effectiveness and robustness in detecting different types of pretraining data in LLMs. We evaluate our method on various sizes of the Pythia model and compare its effectiveness against baseline attacks on seven types of text in the Pile and MIMIR datasets. Table 1 summarizes the results obtained when targeting five Pythia model sizes ranging from 160M to 12B parameters, tested with two configurations of tagged keywords: 4 and 10.
The main findings from these results are as follows:
@@
  • When averaging performance across all datasets, Tag&Tab ($K=4$) outperforms all baseline methods over all the models, establishing itself as the most effective approach overall. Tag&Tab ($K=10$) ranks as the second-best method, demonstrating strong performance but falling short of the results achieved with $K=4$.
  • Notably, Tag&Tab ($K=4$) achieves either the best or second-best results for the majority of textual data types examined. Even when it does not come in first or second place, its performance remains competitive, serving as a robust alternative to the leading method.
@@
  • While Tag&Tab ($K=10$) selects more keywords, increasing the number of probabilities considered for text membership inference, its results are consistently lower than those of Tag&Tab when $K=4$. This supports the hypothesis that selecting a smaller number of keywords allows the method to extract noise-free information from the model.
We also observe that certain formal texts, such as mathematical proofs, may contain fewer named entities or conventional keywords. However, they often feature domain-specific terminology or symbolic expressions that carry strong membership signals. This is evidenced by our results on the DM Mathematics subset (Table 1), where Tag&Tab maintained SOTA performance, achieving an AUC between 0.95 and 0.96.
@@
Despite outperforming the baseline MIAs, the AUC achieved by Tag&Tab is relatively low in certain cases, hovering around 0.55. However, recent research by Maini et al. [17] showed that aggregating multiple MIAs improves dataset membership inference accuracy, emphasizing the need for better attacks that achieve AUC over 0.5 for improved aggregated attack performance. Tag&Tab meets this criterion, making it a valuable component in an ensemble of MIAs, as we further demonstrate in Appendix A.6.
AUC T@F5 AUC T@F5 AUC T@F5 AUC T@F5 AUC T@F5 AUC T@F5 AUC T@F5
Neighbor 0.65 0.27 0.71 0.38 0.90 0.73 0.65 0.26 0.71 0.36 0.96 0.88 0.76 0.48
Loss 0.59 0.25 0.70 0.43 0.89 0.74 0.62 0.24 0.69 0.32 0.97 0.90 0.74 0.48
Zlib 0.53 0.22 0.67 0.42 0.89 0.74 0.55 0.19 0.61 0.25 0.96 0.88 0.70 0.45
Min-20.0% Prob 0.61 0.24 0.70 0.42 0.87 0.70 0.65 0.25 0.70 0.34 0.95 0.86 0.75 0.47
Min‑K%++ (20%) 0.60 0.23 0.68 0.38 0.78 0.60 0.59 0.20 0.56 0.20 0.95 0.86 0.69 0.41
Max-20.0% Prob 0.51 0.15 0.66 0.34 0.87 0.69 0.51 0.13 0.59 0.20 0.96 0.91 0.68 0.40
ReCaLL 0.58 0.22 0.70 0.42 0.84 0.64 0.66 0.29 0.72 0.37 0.74 0.50 0.70 0.41
DC-PDD 0.61 0.27 0.71 0.47 0.88 0.77 0.68 0.34 0.74 0.44 0.95 0.89 0.76 0.53
Ours (Tag&Tab K=4) 0.69 0.28 0.78 0.48 0.91 0.76 0.72 0.30 0.75 0.36 0.97 0.90 0.80 0.51
Ours (Tag&Tab K=10) 0.67 0.26 0.77 0.46 0.91 0.77 0.72 0.30 0.76 0.36 0.96 0.87 0.80 0.50
Table 2: Detection of data from the BookMIA dataset used in pretraining seven models using Tag&Tab and six baseline MIAs, evaluated in terms of AUC and T@F=5%. All results are reported as decimal fractions. The last two rows compare the Tag&Tab method when selecting four and ten keywords. The best results are bolded, and second-best are underscored.
Figure 2: AUC scores as a function of the number of tagged keywords for the examined models on the BookMIA dataset. Yellow points indicate optimal performance: 2-3 keywords for LLaMA 1, 7 for Pythia and GPT-3.5 Turbo, and 4 on average (Average).
5.2 Case Study 2: Detecting Specific Pretraining Data in LLMs
This case study focuses on the precision in detecting specific pretraining data in LLMs. We designed a targeted attack to infer whether copyrighted data was part of the model’s pretraining. Unlike Case Study 1, the member and non-member data come from different sources. Using the BookMIA dataset, we simulate partial knowledge of a model’s pretraining data to infer specific text files suspected of being included in the target model’s training set. For validation, we selected non-members from books published after the target model’s release thus ensuring they were not part of its pretraining data.
To determine the optimal number of keywords to select, we evaluate the results by selecting between 1 and 10 keywords from each sentence The results shown in Figure 2 demonstrate that the optimal number of keywords required to ensure effective detection depends on the model architecture. As can be seen, for different sizes of the LLaMA 1 model, the optimal number of keywords ranged from 2 to 3, while for the Pythia models and GPT-3.5 Turbo, the optimal number of tagged keywords was 7. The best results across all models were achieved with $K=4$, yielding an average AUC score of 0.8.
The main findings from these results are as follows:
  • Tag&Tab outperforms all other attacks in terms of the AUC, with an average improvement of 5.3%–17.6% over baseline methods when $K=4$.
  • Tag&Tab ($K=4$) outperforms other attacks in terms of the T@F=5% on the LLaMA-7B and LLaMA-13B models. However, for most other models, the DC-PDD attack performs better, with Tag&Tab consistently ranking second.
  • As the size of the examined LLM increases, the AUC scores of the MIAs also increase due to the model’s memorization capacity [24]. This can be seen in the results for Tag&Tab which achieved very high AUC scores: (1) 0.91 on LLaMA-30b compared to 0.69 on LLaMA-7b, (2) 0.75 on Pythia-12b compared to 0.72 on Pythia-6.9b, and (3) 0.97 on GPT-3.5 Turbo, the largest model tested.
To better understand our method’s performance, in Appendix A.1, we examine the impact of our method’s tagging stage by comparing the selection of the lowest $K$ entropy words with random token selection, observing that prioritizing the lowest $K$ entropy words significantly enhances performance across all models, resulting in superior AUC scores.
Finally, Appendix A.2 reports the results obtained on additional experiments on our tagging choices. Using only named‑entity tokens or only low‑entropy tokens improves on the baselines, but combining the two as done in Tag&Tab produces the best results. We also experimented with ranking tokens based on TF-IDF rather than entropy, which resulted in noticeably poorer performance.
@@ 6 Conclusion
We present Tag&Tab, a novel black-box method for detecting pretraining data in LLMs. By focusing on the semantic and contextual relevance of words, the method enhances detection capabilities. Tag&Tab outperforms SOTA attacks and consistently achieves high performance across diverse textual data types. Our comprehensive evaluation spans nine textual data types from four datasets (the Pile, MIMIR, BookMIA, and PatentMIA) and ten LLMs of varying sizes and architectures (LLaMA 1-7b, 13b, 30b, Pythia-160m, 1.4b, 2.8b, 6.9b, 12b, GPT-3.5 Turbo, Qwen1.5-14b). Our study confirms that the selection of low-entropy contribution keywords, augmented with NER spans, improves membership inference attack results, further validating our approach. Future work could extend Tag&Tab by considering keyword context and placement within documents. Additionally, developing new MIAs that leverage advanced NLP techniques to assess word significance could further improve the detection of pretraining data in LLMs.
Appendix A Appendix
Our code is available in our GitHub repository.  [^8] https://github.com/sagivantebi/Tag-Tab
A.1 Impact of Our Word Selection Process
To evaluate the impact of our word selection process, Tagging, we compared it against a random selection of words using the same Tabbing algorithm. Tagging selects the lowest $K$ entropy words, augmented with NER spans, while the random baseline selects words without regard to entropy. Figure 3 illustrates the results of this comparison. The dataset used in this evaluation was BookMIA. For each model, we report the AUC scores obtained when selecting between one and ten keywords per text using the Tag&Tab entropy-based method (blue) and a random word selection (orange). The results show that Tag&Tab’s original Tagging process improves performance across all models. Tag&Tab’s original Tagging process achieved an average AUC of 0.8, compared to the average AUC of 0.64 obtained by randomly selecting $K$ words. This demonstrates the Tag method’s effectiveness in enhancing model performance by focusing on rare words.
Figure 3: Comparison of Tag&Tab’s Tabbing performance with our original Tagging keyword selection against Tabbing performance with random word selection for different $K$ values. For each model, we report the AUC scores obtained when selecting 1-10 keywords per text using the Tag&Tab entropy-based with NER spans process (blue) and a random word selection method (orange). Each bar represents the AUC achieved for a specific $K$ value. Results are presented for all evaluated models on the BookMIA dataset as well as the overall average (AVG). The plot highlights the consistent advantage of low-entropy keyword selection across varying $K$ values and model sizes.
A.2 Ablation Study: Alternative Word Selection Strategies
To analyze the contribution of different word selection strategies, we conducted an ablation study comparing several variants of our method on the BookMIA dataset. Specifically, we evaluated the following alternatives to the Tagging process used by the Tag&Tab method:
  • Entropy-Only: selecting the four lowest entropy tokens per sentence.
  • NER-Only: selecting the first four named entities identified via spaCy per sentence.
  • TF-IDF: selecting the top four tokens with the highest TF-IDF scores per sentence.
We evaluated these variants on all of the models except GPT-3.5 Turbo.
The results, summarized in Table 3, show that both the Entropy-Only and NER-Only variants achieved competitive performance, with AUC values slightly below those achieved when the original Tagging process is used by the Tag&Tab method. The TF-IDF variant performed notably worse, confirming that low-entropy and named-entity tokens are more effective indicators of membership when used together. These findings validate our design choice of combining both strategies for optimal pretraining data detection.
Table 3: AUC scores of Tag&Tab and the evaluated variants on the BookMIA dataset, all with $K=4$ selected keywords.
Method LLaMA-7B LLaMA-13B LLaMA-30B Pythia-6.9B Pythia-12B
Tag&Tab 0.69 0.78 0.91 0.72 0.75
Entropy-Only 0.67 0.76 0.89 0.70 0.73
NER-Only 0.62 0.72 0.85 0.68 0.73
TF-IDF 0.60 0.68 0.81 0.65 0.67
Figure 4: Cross-domain membership inference AUC scores using XGBoost classifiers (threshold’s transferability). Each row corresponds to a classifier trained on the domain listed, and each column indicates the evaluation domain. Diagonal values represent in-domain evaluation, while off-diagonal entries show cross-domain transfer performance.
A.3 Threshold Calibration
A.3.1 Calibrating $\gamma$ with No Member Labels
While in a realistic “zero-knowledge” setting, we assume no access to the model’s training data, we can still collect data that were not seen during training. This includes texts published after the model’s release or synthetically generated samples. To calibrate the decision threshold $\gamma$, we run Tag&Tab on a set of known non‑member samples and select $\gamma$ based on the upper tail of Tag&Tab’s scores distribution (e.g., the 95th percentile). This threshold is then used to flag samples that appear “member-like” relative to the known non‑members. Using BookMIA with LLaMA‑30B and K=4K{=}4, this approach yielded $\gamma\approx 0.0392$, resulting in an AUC of 0.85 on the test split, a modest drop from the 0.91 AUC achieved when member labels were available. Although the performance was slightly poorer, the method still achieved strong results, demonstrating its potential for practical real-world applications even under limited knowledge conditions.
A.3.2 Domain Mismatch Test
To examine the threshold’s transferability, we applied the same $\gamma=0.0392$ to a math corpus from MIMIR (DM Mathematics), using the same model, LLaMA‑30B. Performance dropped sharply (AUC <0.60<0.60), showing that Tag&Tab’s scores distribution differs across domains. We saw the same effect when switching models and concluded that each model–dataset pair needs its own calibrated threshold.
A.3.3 Cross-Domain Generalization
To further investigate domain transferability, we extended our thresholding approach beyond a fixed scalar $\gamma$. Instead of using a single cutoff, we trained a supervised XGBoost classifier on Tag&Tab scores extracted from each of seven domains (DM Mathematics, GitHub, Pile CC, C4, Ubuntu IRC, Gutenberg, and EuroParl). In this setup, the XGBoost model functions as a learned, domain-specific thresholding mechanism: it establishes a more flexible decision boundary in score space that generalizes the role of $\gamma$.
Each trained classifier was then evaluated across all domains, yielding the 7×7 AUC matrix shown in Figure 4. This experiment therefore examines whether thresholds calibrated in one domain can transfer effectively to other domains.
The results reveal a clear trend: classifiers achieve the highest AUC when evaluated on the same domain they were trained on (diagonal entries), but their performance drops sharply when applied to different domains. For example, the classifier trained on Ubuntu IRC data achieves an AUC of 0.967 when tested on Ubuntu IRC data but falls below 0.50 on most other domains. Similarly, classifiers trained on Gutenberg (0.990) and EuroParl (0.980) achieve near-perfect within-domain performance but transfer poorly elsewhere.
These findings confirm that Tag&Tab’s inference thresholds are highly domain-specific.
A.3.4 Recommendation
Thresholds should be recalibrated whenever the target model or data domain changes. A small, trustworthy non‑member dev set from the intended domain is sufficient; no labeled members are required.
A.4 Generalization to Chinese Texts
To examine whether Tag&Tab can generalize to a non-Latin language with a fundamentally different structure, we evaluated it on a Chinese text using the PatentMIA dataset [33], which contains patents sourced from Google Patents [43]. This evaluation was performed using the Qwen1.5-14B model, [^9] https://huggingface.co/Qwen/Qwen1.5-14B an open-source LLM developed by Alibaba Cloud, optimized for Chinese and multilingual understanding [39]. Tag&Tab still achieved a good score (AUC = 0.6), but it did not surpass the strongest baseline (DC‑PDD, 0.69 AUC). We attribute this gap to structural differences between Chinese and English, such as the absence of explicit word boundaries and different entropy distributions, which diminish the effectiveness of our current keyword–entropy heuristic.
A.5 Robustness to Adversarial and Distributional Perturbations
To evaluate Tag&Tab’s robustness against minor textual modifications, we conducted an experiment where 2–5 words per sample in the BookMIA dataset were replaced with suitable synonyms. These changes preserved the original meaning while altering the lexical form, simulating both adversarial-style perturbations and natural distribution shifts.
We evaluated the impact of these modifications on the LLaMA-7B, LLaMA-13B, LLaMA-30B, Pythia-6.9B, and Pythia-12B models. In all cases, we used Tag&Tab with $K=4$ keywords.
The results, presented in Table 4, show that Tag&Tab exhibits only a minor performance drop of 1–2% in the AUC across all models. These findings confirm that the method remains effective even when the exact data distribution is unknown, demonstrating resilience to small-scale semantic-preserving shifts.
Table 4: Robustness of Tag&Tab (K=4) under synonym-based perturbations on the BookMIA dataset.
Model Original AUC Perturbed AUC
LLaMA-7B 0.69 0.68
LLaMA-13B 0.78 0.76
LLaMA-30B 0.91 0.89
Pythia-6.9B 0.72 0.71
Pythia-12B 0.75 0.73
A.6 Ensemble Membership Inference Attacks
While Tag&Tab achieves strong standalone results, recent work has shown that ensemble methods, combining multiple MIA scores, often outperform individual attacks [17]. To evaluate whether Tag&Tab score contributes complementary information, we trained a simple XGBoost classifier on features extracted from different attack methods. Experiments were conducted on the BookMIA dataset using the LLaMA-30b model. The feature sets used to train the classifiers included: (i) Tag&Tab scores only, (ii) an ensemble of existing methods’ scores (PPL, Zlib, Min-20%), and (iii) the ensemble augmented with Tag&Tab scores.
The results, summarized in Table 5, demonstrate that Tag&Tab substantially improves ensemble performance. While the ensemble of other methods achieved an AUC of 0.94, adding Tag&Tab raised the AUC to 0.97. These findings confirm that Tag&Tab captures complementary signals not fully exploited by existing approaches, strengthening its value as a component in broader attack pipelines.
Table 5: Ensemble membership inference attack performance on the BookMIA dataset and LLaMA-30b.
Feature Set AUC
Tag&Tab only 0.91
Ensemble (PPL, Zlib, Min-20%) 0.94
Ensemble + Tag&Tab 0.97